Patton is trying to decide between a cup of coffee and a glass of orange juice. The coffee costs $1 and yields 10 units of additional utility. The juice costs $2 and yields 22 units of additional utility. If Patton follows the principle of rational choice, he will buy:

  

juice since it's better for you than coffee

coffee since it gives her the most marginal utility per dollar

juice since it gives her the most marginal utility per dollar

coffee since it is cheaper
Answer #1

Answer
Option 3
juice since it gives her the most marginal utility per dollar


The decision depends on which provides higher per dollar marginal utility
per dollar marginal utility =marginal utility/price

Per dollar marginal utility from a cup of coffee=10/1=10
Per dollar marginal utility from a glass of juice =22/2=11
so the consumer should purchase a glass of juice if he/she has the ability to pay